The private marginal benefit for commodity X is given by 10-X, where Z is the number of units consumed. The private marginal cost of producing X is constant at $5. For each unit of X produced, an external cost of $2 is imposed on member of society.

a. In the absence of any government intervention, how much X is produced?

b. What is the efficient level of production of X?

c. What is the gain to society involved in moving from the inefficient to the efficient level of production?

d. Suggest a Pigouvian tax that would lead to the efficient level. How much revenue would the tax raise?


1
Expert's answer
2021-10-03T18:47:08-0400

a) Private Marginal Benefit (PMB) is a demand curve.

"PMB = 10-X"

Private Marginal Cost (PMC) = 5

PMB = PMC

"10 - X = 5
\\ X = 10 - 5\\ X = 5"

= 5 units

b)

At socially efficient equilibrium, PMB is equal to Social Marginal Cost (SMC).

"= 5 + 2\\
= 7"

Efficient level of production of X will be "10 - 7 = 3"

= 3 units

c)

Gain to society is equal to deadweight loss

"=\frac{1}{2} \times(5-3) \times (7-5)\\
= \frac{1}{2} \times 2 \times 2 \\
= 2"

d)

At socially efficient equilibrium marginal external cost is equal to Pigouvian tax

Revenue is equal to socially equilibrium product multiply by tax.

"= 2\times3\\
= \$6"